OpenStudy (anonymous):

use the formula d=rt to find the value of the missing variable d=57feet, t=19seconds

OpenStudy (anonymous):

since, d=rt i.e. 57=r*19 i.e. r= 57/19 i.e. r= 3 feet/sec

OpenStudy (anonymous):

\[r=\frac{d}{t}\]

OpenStudy (anonymous):

Теперь решить уравнение.